### v1

@@ 4. $(\mathcal{D}_{4})$
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. The order $\prec$ on $M(\Delta_{3}(W_{n}))$ is defined as follows: We say $F\prec F^{\prime}$ if and only if any one of the following conditions is true:
@@
If none of these conditions (i)-(iv) is true, we say $F^{\prime}\prec F$.
@@
Let $F,F^{\prime}\in M(\Delta_{3}(W_{n}))$. Let $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$ for some $s\neq t$ such that ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and ${F^{\prime}}^{c}=({F}^{c}\setminus\{\nu\})\sqcup\{\lambda\}$ for some $\nu\in{F}^{c}$ and $\lambda\in F$. If $\lambda<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$.
@@
Suppose $\nu=\alpha_{s}$. By Remark 3.2, $\nu<_{\mathcal{O}}s_{1},s_{2}$. Since $\lambda<_{\mathcal{O}}\nu$, $\lambda<_{\mathcal{O}}s_{1},s_{2}$. Thus $F^{\prime}\in\mathcal{T}_{t}$ implies that $F^{\prime c}=\{\alpha_{t}\}\sqcup\{s_{1},s_{2}\}$, where $\alpha_{t}=\lambda$ and we get $t<s$ by the definition of $<_{\mathcal{O}}$. Therefore, $F^{\prime}\prec F$ by Definition 3.4.
Now suppose $\nu\in\{s_{1},s_{2}\}$. First, let $\nu=s_{1}$. Then $F^{\prime c}=\{\lambda,\alpha_{s},s_{2}\}$. Since $\lambda<_{\mathcal{O}}\alpha_{s}$ and $F^{\prime c}\in\mathcal{T}_{t}$, either ${F^{\prime}}^{c}=\{\alpha_{t}\}\sqcup\{\alpha_{s},s_{2}\}$ or ${F^{\prime}}^{c}=\{\alpha_{t}\}\sqcup\{s_{2},\alpha_{s}\}$, where $\alpha_{t}=\lambda$. Therefore $t<s$ by the definition of $<_{\mathcal{O}}$, and hence, $F^{\prime}\prec F$ using Definition 3.4. If $\nu=s_{2}$, we get $F^{\prime}\prec F$ using a similar argument. ∎
To prove Theorem 1.4, in Section 3.1, we first show that the order $\prec$ given in Definition 3.4 provides a shelling order for $\Delta_{3}(W_{n})$. Later, in Section 3.2, we identify and count all the spanning facets for this shelling order.
@@ 3.1. Shelling Order
In this section, we prove that the order $\prec$ given in Definition 3.4 provides a shelling order for $\Delta_{3}(W_{n})$. To demonstrate that $\prec$ provides a shelling order for $\Delta_{3}(W_{n})$, by definition of shellability, we need to prove the following: For any $F_{i},F_{j}\in M(\Delta_{3}(W_{n}))$ with $F_{i}\prec F_{j}$, there is a facet $F_{r}$ such that
@@
Let $F_{i}\prec F_{j}$ be two elements of $M(\Delta_{3}(W_{n}))$. If there is an $F\in M(\Delta_{3}(W_{n}))$ such that $F\prec F_{j}$ and $F^{c}=(F_{j}^{c}\setminus\{\nu\})\sqcup\{\lambda\}$ for some $\nu\in F_{j}^{c}$ and $\lambda\in F_{j}\setminus F_{i}$, then $F_{r}=F$ satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
Let $F_{i},F_{j}\in M(\Delta_{3}(W_{n}))$ such that $F_{i}\prec F_{j}$. Let $F_{i}\in\mathcal{T}_{i_{0}}$ and $F_{j}\in\mathcal{T}_{j_{0}}$, where ${F_{i}}^{c}=\{\alpha_{i_{0}}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j_{0}}\}\sqcup\{j_{1},j_{2}\}$. By our assumption, we have $i_{1}<i_{2}$ and $j_{1}<j_{2}$. For convenience, we drop the subscripts $i_{0}$ and $j_{0}$ from $\alpha_{i_{0}}$ and $\alpha_{j_{0}}$, and refer to them simply as $\alpha_{i}$ and $\alpha_{j}$, respectively. Our aim is to find an $F_{r}$ which satisfies ($\ast$3.1) for the pair $F_{i}\prec F_{j}$.
We deal with the cases $F_{j}\in\mathcal{D}$ and $F_{j}\notin\mathcal{D}$ separately. The case $F_{j}\in\mathcal{D}$ is considered in Lemma 3.9, and the case $F_{j}\notin\mathcal{D}$ is considered in Lemmas 3.7 and 3.8. In Lemma 3.7, we deal with the case $i_{0}=j_{0}$, and in Lemma 3.8, we deal with the case $i_{0}\neq j_{0}$.
@@
Suppose that $F_{j}\notin\mathcal{D}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then there exists an $F_{r}\in M(\Delta_{3}(W_{n}))$ that satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
Since $F_{i}\in\mathcal{T}_{j_{0}}$, we have $\alpha_{i}=\alpha_{j}$, and thus ${F_{i}}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$. If $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}\neq\emptyset$, then ($\ast$3.1) is satisfied by taking $F_{r}=F_{i}$. So, we assume that $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$.
Since $F_{i}\prec F_{j}$, $F_{j}\notin\mathcal{D}$ and $F_{i},F_{j}\in\mathcal{T}_{j_{0}}$, using Definition 3.4, we get $F_{i}\notin\mathcal{D}$ and $F_{i}\ll F_{j}$. Further, since $F_{i},F_{j}\in\mathcal{T}_{j_{0}}$, $F_{i}\ll F_{j}$ implies that $F_{i}<_{\mathcal{T}_{j_{0}}}F_{j}$. This means that either $i_{1}<j_{1}$, or $i_{1}=j_{1}$ and $i_{2}<j_{2}$. Since $i_{1},i_{2},j_{1}$, and $j_{2}$ are distinct, we get $i_{1}<j_{1}<j_{2}$. Moreover, $i_{1}<i_{2}$. By Remark 3.3, we have $22\leq\alpha_{j}\leq n-$ 2.
@@
    Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Since $i_{1}<j_{1}<j_{2}<\alpha_{j}$, we have $i_{1}<\alpha_{j}-2$. Suppose $i_{1}\sim\alpha_{j}$. Then $\alpha_{j}\leq n-2$ implies that $i_{1}=0$ and $\alpha_{j}=n-2$. Since $W_{n}[{F_{i}}^{c}]$ is disconnected, we get $i_{2}\neq n-1$ Therefore, $i_{1},i_{2},j_{1},j_{2}<\alpha_{j}$. We have $\alpha_{j}=n-2>m$ as $n\geq 9$) and $\alpha_{j}<_{\mathcal{O}}i_{1},i_{2},j_{1},j_{2}$. Thus, by Remark 3.1 (ii), $i_{1},i_{2},j_{1},j_{2}<2m-\alpha_{j}=2m-n+2\leq 3$, a contradiction as $i_{1},i_{2},j_{1},j_{2}$ are distinct. Hence $i_{1}\nsim\alpha_{j}.$ Since $W_{n}[{F_{j}}^{c}]$ is disconnected, we have $j_{2}\nsim\alpha_{j}$ or $j_{2}\nsim j_{1}$. If $j_{2}\nsim\alpha_{j}$, then $\alpha_{j}$ is an isolated vertex in $W_{n}[F^{\prime c}]$; and if $j_{2}\nsim j_{1}$, then $j_{2}\nsim i_{1}$, which implies that $i_{1}$ is an isolated vertex in $W_{n}[F^{\prime c}]$. This means that $W_{n}[{F^{\prime}}^{c}]$ is disconnected and hence $F^{\prime}\in M\Delta_{3}(W_{n}))$.
    Since $i_{1}<j_{1}<j_{2}<\alpha_{j}$, we observe that $F^{\prime}$ does not satisfy any of the conditions $(\mathcal{D}_{1})$-$(\mathcal{D}_{4})$, and therefore $F^{\prime}\not\in\mathcal{D}$. We have $F^{\prime}\in\mathcal{T}_{j_{0}}$. So, $i_{1}<j_{1}$ implies that $F^{\prime}<_{\mathcal{T}_{j_{0}}}F$ and thus $F^{\prime}\ll F_{j}$. Hence $F^{\prime}\prec F_{j}$ by Definition 3.4 (i), and thus, $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
@@
    Since $W_{n}[{F_{j}}^{c}]$ is disconnected, we get $j_{2}>\alpha_{j}+3$. Thus, $\alpha_{j}\geq 2$ implies that $j_{2}\nsim\alpha_{j}$. Further, $F_{i}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},i_{2}\}$. So we have either $i_{1}<\alpha_{j}$, or $i_{1}>\alpha_{j}$.
@@
      Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$. We show that if $F^{\prime}\in M\Delta_{3}(W_{n}))$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1), otherwise $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). First, assume that $F^{\prime}\in M\Delta_{3}(W_{n}))$. Using the facts that $i_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}+3$, we see that $F^{\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$. Therefore $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
      We now assume that $F^{\prime}\notin M\Delta_{3}(W_{n}))$ i.e., $W_{n}[F^{\prime c}]$ is connected. Since $j_{2}\nsim\alpha_{j}$, we get $\alpha_{j}\sim i_{1}$ and $i_{1}\sim j_{2}.$ Thus, $j_{2}>j_{1}>\alpha_{j}>i_{1}$ implies that $i_{1}\geq\alpha_{j}-2$. We know that $i_{1}<i_{2}$ and $W_{n}[F_{i}^{c}]$ is disconnected. So $i_{2}>\alpha_{j}$. Since $i_{1},i_{2},j_{1}$ and $j_{2}$ are distinct, it follows that $i_{2}\leq n-3$ or $j_{1}\leq n-3$.
      We have $i_{2},j_{1}>\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}i_{2},j_{1}$. If $\alpha_{j}<m$, then using Remark 3.1 (i), $i_{2},j_{1}\geq 2m-\alpha_{j}$; and if $\alpha_{j}>m$, then $i_{2},j_{1}>\alpha_{j}>2m-\alpha_{j}$. Hence we conclude that $i_{2},j_{1}\geq 2m-\alpha_{j}$. This means that $n-3\geq 2m-\alpha_{j}\geq n-\alpha_{j}$. Hence $\alpha_{j}\geq 3$. Now $i_{1}\geq\alpha_{j}-2$ and $i_{1}\sim j_{2}$ imply that $\alpha_{j}=3$, $i_{1}=1$ and $j_{2}=n-1.$ If $2m=n+1,$ then $2m-\alpha_{j}=n-2>n-3,$ a contradiction. So $2m=n.$ Then $32m-\alpha_{j}=n-$ 3, and this implies that $j_{1}\in\{n-3,n-2\}$. Hence $j_{1}\geq 6$ (as $n\geq 9$). Therefore, since $i_{1}=1$ and $\alpha_{j}=3$, it follows that $j_{1}\nsim i_{1}$ and $j_{1}\nsim\alpha_{j}$. Clearly, $W_{n}[F^{\prime\prime c}]$ is disconnected i.e., $F^{\prime\prime}\in M\Delta_{3}(W_{n}))$. Now, using the facts that $i_{1}=\alpha_{j}-2$ and $j_{1}\geq\alpha_{j}+3$, we have $F^{\prime\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime\prime}\ll F_{j}$. Thus $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (i). Hence $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.6.
@@
      Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Suppose $W_{n}[{F^{\prime}}^{c}]$ is connected. Then $2\leq\alpha_{j}<i_{1}<j_{1}<j_{2}$ implies that $i_{1}\leq\alpha_{j}+2$, $j_{1}=i_{1}+1$ and $j_{2}=i_{1}+2$. If $i_{1}=\alpha_{j}+1$, then $j_{1}=\alpha_{j}+2$ and $j_{2}=\alpha_{j}+3$, which implies that $W_{n}[{F_{j}}^{c}]$ is connected, a contradiction. So $i_{1}=\alpha_{j}+2$. This means that $j_{1}=\alpha_{j}+3$ and $j_{2}=\alpha_{j}+4$. Since $i_{1}<i_{2}$ and $i_{1},i_{2},j_{1},j_{2}$ are distinct, $\alpha_{j}\leq n-6$. Further, by Remark 3.1 (i) and (ii), $i_{1}>\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}i_{1}$ implies that $i_{1}\geq 2m-\alpha_{j}$ i.e., $\alpha_{j}+2\geq 2m-\alpha_{j}$. Hence, $\alpha_{j}\geq m-1$. This implies that $F_{j}$ satisfies $(\mathcal{D}_{4})$, which contradicts our assumption that $F_{j}\notin\mathcal{D}$. Therefore $W_{n}[{F^{\prime}}^{c}]$ is disconnected and hence $F^{\prime}\in M\Delta_{3}(W_{n}))$.
      Since $\alpha_{j}<i_{1}<j_{1}<j_{2}$, we have $F^{\prime}\notin\mathcal{D}$. Further, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, and thus $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
@@
    Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$, ${F^{\prime\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$ and $F^{\prime\prime\prime c}=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$.
@@
    First, let $F^{\prime\prime}\in M\Delta_{3}(W_{n}))$. Since $i_{1}<j_{1}$, we have $F^{\prime\prime}\ll F$. Therefore, if $F^{\prime\prime}\notin\mathcal{D}$, then $F^{\prime\prime}\prec F_{j}$ by Definition 3.4 (i). Hence, by taking $F_{r}=F^{\prime\prime}$, ($\ast$3.1) is satisfied.
@@
    Suppose $F^{\prime\prime}\in\mathcal{D}$. Since $i_{1}<j_{1}<\alpha_{j}$, it follows that $F^{\prime\prime}$ satisfies $(\mathcal{D}_{3})$. This means that $i_{1}=\alpha_{j}-4$, $j_{1}=\alpha_{j}-3$, and $\alpha_{j}=m+1$. Note that $n\geq 9$ implies $m\geq 5$. So $\alpha_{j}\geq 6$. Now, $i_{1}=\alpha_{j}-4\geq 2$ and $j_{2}>\alpha_{j}$ implies that $i_{1}\nsim\alpha_{j}$ and $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(W_{n}))$ and $F^{\prime}\notin\mathcal{D}$. Also, since $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, $F^{\prime}\prec F_{j}$ by Definition 3.4 (i). Therefore, taking $F_{r}=F^{\prime}$, ($\ast$3.1) is satisfied by Proposition 3.6.

### v2

@@ 4. $(\mathcal{D}_{4})$
Let $F,F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. The order $\prec$ on $M(\Delta_{3}(C_{n}^{2}))$ is defined as follows: We say $F\prec F^{\prime}$ if and only if any one of the following conditions is true:
@@
It can be easily observed that $\prec$ is a total order.
@@
Let $F,F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Let $F\in\mathcal{T}_{s}$ and $F^{\prime}\in\mathcal{T}_{t}$ for some $s\neq t$ such that ${F}^{c}=\{\alpha_{s}\}\sqcup\{s_{1},s_{2}\}$ and ${F^{\prime}}^{c}=({F}^{c}\setminus\{\nu\})\sqcup\{\lambda\}$ for some $\nu\in{F}^{c}$ and $\lambda\in F$. If $\lambda<_{\mathcal{O}}\alpha_{s}$, then $F^{\prime}\prec F$.
@@
Since $\alpha_{s}<_{\mathcal{O}}s_{1},s_{2}$ by Remark 3.2, $\lambda<_{\mathcal{O}}\alpha_{s}$ implies that $\lambda<_{\mathcal{O}}s_{1},s_{2}$. Therefore $F^{\prime c}=\{\alpha_{t}\}\sqcup\{t_{1},t_{2}\}$, where $\alpha_{t}=\lambda$ such that $t<s$ and $t_{1},t_{2}\in F^{c}$. Hence $F^{\prime}\prec F$ by Definition 3.5. ∎
To prove Theorem 1.4, in Section 3.1, we first show that the order $\prec$ given in Definition 3.5 provides a shelling order for $\Delta_{3}(C_{n}^{2})$. Later, in Section 3.2, we identify and count all the spanning facets for this shelling order.
@@ 3.1. Shelling Order
In this section, we prove that the order $\prec$ given in Definition 3.5 provides a shelling order for $\Delta_{3}(C_{n}^{2})$. To demonstrate that $\prec$ provides a shelling order for $\Delta_{3}(C_{n}^{2})$, by definition of shellability, we need to prove the following: For any $F_{i},F_{j}\in M(\Delta_{3}(C_{n}^{2}))$ with $F_{i}\prec F_{j}$, there is a facet $F_{r}$ such that
@@
Let $F_{i}\prec F_{j}$ be two elements of $M(\Delta_{3}(C_{n}^{2}))$. If there is an $F\in M(\Delta_{3}(C_{n}^{2}))$ such that $F\prec F_{j}$ and $F^{c}=(F_{j}^{c}\setminus\{\nu\})\sqcup\{\lambda\}$ for some $\nu\in F_{j}^{c}$ and $\lambda\in F_{j}\setminus F_{i}$, then $F_{r}=F$ satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
Let $F_{i},F_{j}\in M(\Delta_{3}(C_{n}^{2}))$ such that $F_{i}\prec F_{j}$. Let $F_{i}\in\mathcal{T}_{i_{0}}$ and $F_{j}\in\mathcal{T}_{j_{0}}$, where ${F_{i}}^{c}=\{\alpha_{i_{0}}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j_{0}}\}\sqcup\{j_{1},j_{2}\}$. By our assumption, we have $i_{1}<i_{2}$ and $j_{1}<j_{2}$. For convenience, we drop the subscripts $i_{0}$ and $j_{0}$ from $\alpha_{i_{0}}$ and $\alpha_{j_{0}}$, and refer to them simply as $\alpha_{i}$ and $\alpha_{j}$, respectively. Our aim is to find an $F_{r}$ which satisfies ($\ast$3.1) for the pair $F_{i}\prec F_{j}$.
We deal with the cases $F_{j}\in\mathcal{D}$ and $F_{j}\notin\mathcal{D}$ separately. The case $F_{j}\in\mathcal{D}$ is considered in Lemma 3.10, and the case $F_{j}\notin\mathcal{D}$ is considered in Lemmas 3.8 and 3.9. In Lemma 3.8, we deal with the case $i_{0}=j_{0}$, and in Lemma 3.9, we deal with the case $i_{0}\neq j_{0}$. Recall that for any $x,y\in V(C_{n}^{2})$, $x\sim y$ denotes that $x$ is adjacent to $y$ in $C_{n}^{2}$ i.e., $\{x,y\}\in E(C_{n}^{2})$.
@@
Suppose that $F_{j}\notin\mathcal{D}$. If $F_{i}\in\mathcal{T}_{j_{0}}$, then there exists an $F_{r}\in M(\Delta_{3}(C_{n}^{2}))$ that satisfies (italic-$\ast$3.1) for the pair $F_{i}\prec F_{j}$.
@@
Since $F_{i}\in\mathcal{T}_{j_{0}}$, we have $\alpha_{i}=\alpha_{j}$, and thus ${F_{i}}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},i_{2}\}$ and ${F_{j}}^{c}=\{\alpha_{j}\}\sqcup\{j_{1},j_{2}\}$. Therefore, $\alpha_{j}<_{\mathcal{O}}i_{1},i_{2},j_{1},j_{2}$ by Remark 3.2. If $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}\neq\emptyset$, then ($\ast$3.1) is satisfied by taking $F_{r}=F_{i}$. So, we assume that $\{i_{1},i_{2}\}\cap\{j_{1},j_{2}\}=\emptyset$.
Since $F_{i}\prec F_{j}$, $F_{j}\notin\mathcal{D}$ and $F_{i},F_{j}\in\mathcal{T}_{j_{0}}$, using Definition 3.5, we get $F_{i}\notin\mathcal{D}$ and $F_{i}\ll F_{j}$. Therefore $F_{i},F_{j}\in\mathcal{T}_{j_{0}}$ implies that either $i_{1}<j_{1}$, or $i_{1}=j_{1}$ and $i_{2}<j_{2}$ by Definition 3.4 (i). Since $i_{1},i_{2},j_{1}$, and $j_{2}$ are distinct, we get $i_{1}<j_{1}<j_{2}$. Moreover, $i_{1}<i_{2}$. By Remark 3.3, we have $22\leq\alpha_{j}\leq n-$ 2.
@@
    Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Since $i_{1}<j_{1}<j_{2}<\alpha_{j}$, we have $i_{1}<\alpha_{j}-2$. Suppose $i_{1}\sim\alpha_{j}$. Then $\alpha_{j}\leq n-2$ implies that $i_{1}=0$ and $\alpha_{j}=n-2$. Since $C_{n}^{2}[{F_{i}}^{c}]$ is disconnected, we get $i_{2}\neq n-1$. Therefore, $i_{1},i_{2},j_{1},j_{2}<\alpha_{j}$. We have $\alpha_{j}=n-2>m$ (as $n\geq 9$) and $\alpha_{j}<_{\mathcal{O}}i_{1},i_{2},j_{1},j_{2}$. Thus, by Remark 3.1 (ii), $i_{1},i_{2},j_{1},j_{2}<2m-\alpha_{j}=2m-n+2\leq 3$, a contradiction as $i_{1},i_{2},j_{1},j_{2}$ are distinct. Hence $i_{1}\nsim\alpha_{j}.$ Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, we have $j_{2}\nsim\alpha_{j}$ or $j_{2}\nsim j_{1}$. If $j_{2}\nsim\alpha_{j}$, then $\alpha_{j}$ is an isolated vertex in $C_{n}^{2}[F^{\prime c}]$; and if $j_{2}\nsim j_{1}$, then $j_{2}\nsim i_{1}$, which implies that $i_{1}$ is an isolated vertex in $C_{n}^{2}[F^{\prime c}]$. This means that $C_{n}^{2}[{F^{\prime}}^{c}]$ is disconnected, and hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$.
    Since $i_{1}<j_{1}<j_{2}<\alpha_{j}$, we observe that $F^{\prime}$ does not satisfy any of the conditions $(\mathcal{D}_{1})$-$(\mathcal{D}_{4})$, and therefore $F^{\prime}\not\in\mathcal{D}$. We have $F^{\prime}\in\mathcal{T}_{j_{0}}$. So, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$ by Definition 3.4 (i). Hence $F^{\prime}\prec F_{j}$ by Definition 3.5 (i), and thus, $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
@@
    Since $C_{n}^{2}[{F_{j}}^{c}]$ is disconnected, we get $j_{2}>\alpha_{j}+3$. Thus, $\alpha_{j}\geq 2$ implies that $j_{2}\nsim\alpha_{j}$. Further, $F_{i}^{c}=\{\alpha_{j}\}\sqcup\{i_{1},i_{2}\}$. So we have either $i_{1}<\alpha_{j}$, or $i_{1}>\alpha_{j}$.
@@
      Let $F^{\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$ and $F^{\prime\prime c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$. We show that if $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$, then $F_{r}=F^{\prime}$ satisfies ($\ast$3.1), otherwise $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1). First, assume that $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Using the facts that $i_{1}<\alpha_{j}$ and $j_{2}>\alpha_{j}+3$, we see that $F^{\prime}\notin\mathcal{D}$. Also, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$ by Definition 3.4 (i). Therefore, $F^{\prime}\prec F_{j}$ by Definition 3.5 (i). Hence $F_{r}=F^{\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
      We now assume that $F^{\prime}\notin M(\Delta_{3}(C_{n}^{2}))$ i.e., $C_{n}^{2}[F^{\prime c}]$ is connected. Since $j_{2}\nsim\alpha_{j}$, we get $\alpha_{j}\sim i_{1}$ and $i_{1}\sim j_{2}.$ Thus, $j_{2}>j_{1}>\alpha_{j}>i_{1}$ implies that $i_{1}\geq\alpha_{j}-2$. We know that $i_{1}<i_{2}$ and $C_{n}^{2}[F_{i}^{c}]$ is disconnected. So $i_{2}>\alpha_{j}$. Since $i_{1},i_{2},j_{1}$ and $j_{2}$ are distinct, it follows that $i_{2}\leq n-3$ or $j_{1}\leq n-3$.
      We have $i_{2},j_{1}>\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}i_{2},j_{1}$. If $\alpha_{j}<m$, then using Remark 3.1 (i), $i_{2},j_{1}\geq 2m-\alpha_{j}$; and if $\alpha_{j}>m$, then $i_{2},j_{1}>\alpha_{j}>2m-\alpha_{j}$. Hence we conclude that $i_{2},j_{1}\geq 2m-\alpha_{j}$. This means that $n-3\geq 2m-\alpha_{j}\geq n-\alpha_{j}$. Hence $\alpha_{j}\geq 3$. Now $i_{1}\geq\alpha_{j}-2$ and $i_{1}\sim j_{2}$ imply that $\alpha_{j}=3$, $i_{1}=1$ and $j_{2}=n-1.$ If $2m=n+1,$ then $2m-\alpha_{j}=n-2>n-3,$ a contradiction. So $2m=n.$ Then $32m-\alpha_{j}=n-$ 3, and this implies that $j_{1}\in\{n-3,n-2\}$. Hence $j_{1}\geq 6$ (as $n\geq 9$). Therefore, since $i_{1}=1$ and $\alpha_{j}=3$, it follows that $j_{1}\nsim i_{1}$ and $j_{1}\nsim\alpha_{j}$. Clearly, $C_{n}^{2}[F^{\prime\prime c}]$ is disconnected i.e., $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Now, using the facts that $i_{1}=\alpha_{j}-2$ and $j_{1}\geq\alpha_{j}+3$, we have $F^{\prime\prime}\notin\mathcal{D}$. Also $i_{1}<j_{1}$ implies that $F^{\prime\prime}\ll F_{j}$ by Definition 3.4 (i). Thus $F^{\prime\prime}\prec F_{j}$ by Definition 3.5 (i). Hence $F_{r}=F^{\prime\prime}$ satisfies ($\ast$3.1) by Proposition 3.7.
@@
      Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$. Suppose $C_{n}^{2}[{F^{\prime}}^{c}]$ is connected. Then $2\leq\alpha_{j}<i_{1}<j_{1}<j_{2}$ implies that $i_{1}\leq\alpha_{j}+2$, $j_{1}=i_{1}+1$ and $j_{2}=i_{1}+2$. If $i_{1}=\alpha_{j}+1$, then $j_{1}=\alpha_{j}+2$ and $j_{2}=\alpha_{j}+3$, which implies that $C_{n}^{2}[{F_{j}}^{c}]$ is connected, a contradiction. So $i_{1}=\alpha_{j}+2$. This means that $j_{1}=\alpha_{j}+3$ and $j_{2}=\alpha_{j}+4$. Since $i_{1}<i_{2}$ and $i_{1},i_{2},j_{1},j_{2}$ are distinct, $\alpha_{j}\leq n-6$. Further, by Remark 3.1 (i) and (ii), $i_{1}>\alpha_{j}$ and $\alpha_{j}<_{\mathcal{O}}i_{1}$ implies that $i_{1}\geq 2m-\alpha_{j}$ i.e., $\alpha_{j}+2\geq 2m-\alpha_{j}$. Hence, $\alpha_{j}\geq m-1$. This implies that $F_{j}$ satisfies $(\mathcal{D}_{4})$, which contradicts our assumption that $F_{j}\notin\mathcal{D}$. Therefore $C_{n}^{2}[{F^{\prime}}^{c}]$ is disconnected and hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$.
      Since $\alpha_{j}<i_{1}<j_{1}<j_{2}$, we have $F^{\prime}\notin\mathcal{D}$. Further, $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$ by Definition 3.4 (i), and thus $F^{\prime}\prec F_{j}$ by Definition 3.5 (i). Therefore $F_{r}=F^{\prime}$ satisfies ($\ast$3.1).
@@
    Let ${F^{\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{2}\}$, ${F^{\prime\prime}}^{c}:=\{\alpha_{j}\}\sqcup\{i_{1},j_{1}\}$ and $F^{\prime\prime\prime c}:=\{\alpha_{j}\}\sqcup\{j_{1},i_{2}\}$.
@@
    First, let $F^{\prime\prime}\in M(\Delta_{3}(C_{n}^{2}))$. Since $i_{1}<j_{1}$, we have $F^{\prime\prime}\ll F$ by Definition 3.4 (i). Therefore, if $F^{\prime\prime}\notin\mathcal{D}$, then $F^{\prime\prime}\prec F_{j}$ by Definition 3.5 (i). Hence, by taking $F_{r}=F^{\prime\prime}$, ($\ast$3.1) is satisfied.
@@
    Suppose $F^{\prime\prime}\in\mathcal{D}$. Since $i_{1}<j_{1}<\alpha_{j}$, it follows that $F^{\prime\prime}$ satisfies $(\mathcal{D}_{3})$. This means that $i_{1}=\alpha_{j}-4$, $j_{1}=\alpha_{j}-3$, and $\alpha_{j}=m+1$. Note that $n\geq 9$ implies $m\geq 5$. So $\alpha_{j}\geq 6$. Now, $i_{1}=\alpha_{j}-4\geq 2$ and $j_{2}>\alpha_{j}$ implies that $i_{1}\nsim\alpha_{j}$ and $i_{1}\nsim j_{2}$. Hence $F^{\prime}\in M(\Delta_{3}(C_{n}^{2}))$ and $F^{\prime}\notin\mathcal{D}$. Also, since $i_{1}<j_{1}$ implies that $F^{\prime}\ll F_{j}$, we get $F^{\prime}\prec F_{j}$ by Definition 3.5 (i). Therefore, taking $F_{r}=F^{\prime}$, ($\ast$3.1) is satisfied by Proposition 3.7.